OpenStudy (anonymous):

The sales tax in the city of Taxemalot is 8.1%. What is the growth factor for any item purchased at your local Tarmart?

OpenStudy (dan815):

lol taxemalot

OpenStudy (dan815):

lets say the price of something you buy = x so tax = x*8.1/100 = x*0.081 total price you pay is then x+x*0.081= x(1+0.081)=x(1.081) this 1.081 is called the growth factor i think